# Quaternion Multi-focus Color Image Fusion

### v1

@@ III-B2 Optimization
1 Initialize ${\{\dot{\bm{\mathrm{J}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Z}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{0}}\}}{{}_{k=1}^{K}}$,${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{0}}\}}{{}_{k=1}^{K}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}^{0}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}^{0}}$, $\dot{\bm{\mathrm{B}}}{{}^{0}}$, $\dot{\bm{\mathrm{D}}}{{}^{0}}$, $\dot{\bm{\mathrm{E}}}{{}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}^{0}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{5}^{0}}$;
@@ 9      Fix other variables and solve $\dot{\bm{\mathrm{D}}}{{}^{t}}$ using Lemma 2;
12             $\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{t-1}}+\mu(\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}}-\dot{\bm{\mathrm{J}}}{{}_{k}^{t}})$;
13            $\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t}}=\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{t-1}}+\mu(\mathcal{R}(\dot{\bm{\mathrm{B}}}){{}_{k}^{t}}-\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}^{t}})$;
@@
16      ${\dot{\bm{\mathrm{Y}}}{{}_{3}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{3}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{1}^{t}}-\nabla_{1}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
17      ${\dot{\bm{\mathrm{Y}}}{{}_{4}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{4}^{t-1}}}+\mu(\dot{\bm{\mathrm{G}}}{{}_{2}^{t}}-\nabla_{2}\dot{\bm{\mathrm{B}}}{{}^{t}})$;
18      ${\dot{\bm{\mathrm{Y}}}{{}_{5}^{t}}}={\dot{\bm{\mathrm{Y}}}{{}_{5}^{t-1}}}+\mu(\dot{\bm{\mathrm{I}}}-\dot{\bm{\mathrm{B}}}{{}^{t}}-\dot{\bm{\mathrm{D}}}{{}^{t}}-\dot{\bm{\mathrm{E}}}{{}^{t}})$;
@@ 20      $t$longleftarrow $t+1$;
Algorithm 1 Quaternion sparse decomposition
@@ III-C Quaternion base-detail fusion strategy
This subsection proposes a quaternion base-detail fusion (QBDF) strategy to achieve high-quality color image fusion in base and detail scales. Fig. 3 illustrates our QBDF strategy. Given quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{2}$ $\in\mathbbm{H}^{2\times M\times N}$ of two input color images, our QBDF strategy first uses the QSD model to learn their detail-scale layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{2}$ and coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{2}$, and then performs two-scale image fusion.
To fully utilize the complementary features of coefficient matrices and detail-scale layers, QBDF strategy introduces a two-scale fusion process to obtain base-scale and detail-scale fusion results at the patch level. The patch-wise fusion rule is adopted to reduce the sensitivity to the undetermined pixels directly generated by focus detection. First, we define two-scale focus measures as below.
$l_{b,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1}+\theta{\|\dot{\bm{\mathrm{Z}}}_{j}^{(i)}\|}_{2},$ (13)
$l_{d,j}=\mathop{\arg\max}_{i}\ \ \sum_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}_{1},$ (14)
where $l_{b}$ and $l_{d}$ denote the patch-wise focus maps of base-scale and detail-scale fused results respectively, $j$ defines the center spatial location of each patch and parameter $\theta$ is here set to 1 by default. $\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}$ represents the $j^{th}$ patch of $\dot{\bm{\mathrm{D}}}^{(i)}$ and $\dot{\bm{\mathrm{Z}}}{{}^{(i)}_{j}}$ corresponds to the $j^{th}$ column vector of $\dot{\bm{\mathrm{Z}}}^{(i)}$. The detail-scale focus measure uses the two-directional gradients of the detail-scale layers. The gradients that intuitively represent variations in the quaternion pixels can substantially reflect the focus levels of color images. Based on the greater amplitude of Eq. (14), $l_{d,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
Figure 3: Flowchart of our quaternion base-detail fusion (QBDF) and quaternion structural similarity refinement (QSSR) strategies.
Although the detail-scale focus map $l_{d}$ effectively corrects the inaccurate focus map using the neighboring information for boundary and small focused regions, it may produce visual artifacts and fail to handle low-gradient background areas. This limitation can lead to significant inconsistencies in the detail-scale fused result. To address this issue, the base-scale focus measure integrates the gradients of the detail layers with their corresponding coefficient matrices. This ensures local consistency and mitigates spatial artifacts caused by the detail-scale focus map. Based on the greater amplitude of Eq. (13), $l_{b,j}$ is assigned with the corresponding $i$ in $j^{th}$ patch.
The focused patches are obtained and aggregated into detail-scale and the base-scale fused results ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using the following two-scale patch-wise fusion rule:
$\dot{\bm{\mathrm{P}}}{{}^{(1)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{d,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{d,j}=2\end{aligned}\right.$ (15)
$\dot{\bm{\mathrm{P}}}{{}^{(2)}_{j}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{I}}}{{}^{(1)}_{j}},&&l_{b,j}=1\\ &\dot{\bm{\mathrm{I}}}{{}^{(2)}_{j}},&&l_{b,j}=2\end{aligned}\right.$ (16)
@@ III-D Quaternion structural similarity refinement strategy
To obtain a high-quality fused result, we introduce a quaternion structural similarity refinement (QSSR) strategy to refine the base-scale and detail-scale fused patches of ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ using a selection rule with a maximum quaternion image structural similarity value. We assume that the patch to be selected should be most similar to the corresponding region of the source image with the largest focus information. Therefore, we propose a quaternion weighted structural similarity (QWSSIM) strategy to refine the distorted patches that are wrongly detected as focused ones based on the theory of Definition 5. QWSSIM is defined in the following equation.
$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$ (17)
$QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})=\sum_{i=1}^{2}{\tau}_{i,j}QSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)},{\dot{\bm{\mathrm{I}}}}_{j}^{(i)}),$
where $j\in[1,P]$ denotes the center spatial location of each patch where $P$ denotes the number of overlapping patches. ${\tau}_{i,j}$ is an assigned weight to $j^{th}$ quaternion patch of $i^{th}$ input quaternion representation. The patches selected from the two-scale fused results by our QBDF strategy should be quite similar to the corresponding focused patches of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}^{2}_{i=1}$. In other words, each local quaternion patch of the final fused result has a maximum QWSSIM value with the corresponding region of ${\{\dot{\bm{\mathrm{I}}}\}}^{2}_{i=1}$. ${\tau}_{i,j}$ is significantly larger when the corresponding region of ${\dot{\bm{\mathrm{I}}}}^{(i)}$ is focused. Fig. 3 illustrates the flowchart of our QSSR strategy. Our QSSR strategy further compares two-scale fused results with the input quaternion representations patch-by-patch to assess the patch-wise structural similarity using our QWSSIM measure. Patches with higher QWSSIM values are selected to form the final fused result.
To adaptively compute ${\tau}_{i,j}$, we utilize the detail-scale measure to generate each weight map with the size of $n\times P$ where n depends on the number of inputs. The detail-scale measure is enhanced using a Laplacian function $\phi(x)=1-e^{-\frac{x}{\gamma}}$ where $\gamma$ is $=0.2$ by default. Then the weight map is formulated in the following equation.
${\tau}_{i,j}=\frac{\phi({({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}})})}{\sum_{i}\phi(({\sum}_{t=1}^{2}{\|\nabla_{t}\dot{\bm{\mathrm{D}}}{{}^{(i)}_{j}}\|}{{}_{1}}))}.$ (18)
We apply a maximum selection rule based on the QWSSIM measure at the patch level. It effectively corrects defocused regions in the two-scale fused images. Additionally, this strategy eliminates visual artifacts caused by inaccurate focus maps. The selection rule is defined as follows.
${\bm{\mathrm{W}}}_{j}=\left\{\begin{aligned} &1,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})>QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)})\\ &0,&QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(1)})\leq QWSSIM({\dot{\bm{\mathrm{P}}}}_{j}^{(2)}).\end{aligned}\right.$ (19)
The weight matrix $\bm{\mathrm{W}}$ denotes a refined patch-wise focus map. When ${\bm{\mathrm{W}}}_{j}=1$, we select the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(1)}$. Otherwise, we choose the $j^{th}$ patch from ${\dot{\bm{\mathrm{P}}}}^{(2)}$. The final fused result is obtained at the patch level as below:
$\dot{\bm{\mathrm{F}}}_{j}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{P}}}}_{j}^{(1)},&{\bm{\mathrm{W}}}_{j}=1\\ &{\dot{\bm{\mathrm{P}}}}_{j}^{(2)},&{\bm{\mathrm{W}}}_{j}=0.\end{aligned}\right.$ (20)
The patches are aggregated to form a quaternion fused result ${\dot{\bm{\mathrm{F}}}}$. After the inverse transformation from the quaternion domain into the real domain, we obtain a fused color image $\bm{\mathrm{F}}$ $\in\mathbbm{R}^{M\times N\times 3}$.
III-E Extension to multiple color image fusion
Fig. 1 displays our quaternion multi-focus color image fusion framework to fuse two partially focused color images. Our framework is flexible to be extended into fusing multiple color images. Let ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ denote $n$ input color images ($n\geq 2$). The multi-focus color image fusion performed on ${\{{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ is decribed in the following process.
Focus detection. We first encode multiple color images into their quaternion representations ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$. The QSD model is then used to decompose them into quaternion coefficient matrices ${\{\dot{\bm{\mathrm{Z}}}^{(i)}\}}_{i=1}^{n}$ and quaternion detail layers ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ in the iterative learning process.
QBDF strategy. Our QBDF strategy uses two-scale focus measures to obtain the fused results. The detail-scale focus measure uses two-directional gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ while the base-scale focus measure integrates the gradients of $\dot{\bm{\mathrm{D}}}^{(i)}$ with their corresponding $\dot{\bm{\mathrm{Z}}}^{(i)}$ for each quaternion representation. The base-scale and detail-scale measures are easily extended to ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ ($n\geq 2$) using Eqs. (13) and (14). We select focused patches to obtain base-scale and detail-scale fused results ${\{\dot{\bm{\mathrm{P}}}^{(i)}\}}_{i=1}^{2}$ based on Eqs. (15) and (16) similarly.
QSSR strategy. Our QSSR strategy compares each patch from base-scale and detail-scale fused results with input quaternion representations respectively using the QWSSIM. The patches with a higher QWSSIM value are selected as focused ones to form a final fused result. Therefore, the QBDF and QSSR strategies are easily extended to multiple color images. The detailed process is presented in Algorithm 2.
Input: Input $n$ color images ${\{\bm{\mathrm{I}}^{(i)}\}}{{}_{i=1}^{n}}$ ($n\geq 2$)
Output: The fused image $\bm{\mathrm{F}}$
1 for $i=1,\cdots,n$ do
2       Transform $\bm{\mathrm{I}}^{(i)}\in\mathbbm{R}$ into $\dot{\bm{\mathrm{I}}}^{(i)}$ in the quaternion domain;
3      Solve the quaternion sparse coefficient matrix ${\dot{\bm{\mathrm{Z}}}}^{(i)}$ and the quaternion detail layer ${\dot{\bm{\mathrm{D}}}}^{(i)}$ using Algorithm 1;
4 end for
5Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{1}$ overlapping patches respectively;
6Fuse the detail-scale result ${\dot{\bm{\mathrm{P}}}}^{(1)}$ using Eq. (14) ;
7Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P_{2}$ overlapping patches respectively;
8Fuse the base-scale result ${\dot{\bm{\mathrm{P}}}}^{(2)}$ using Eq. (13);
9Divide ${\dot{\bm{\mathrm{P}}}}^{(1)}$ and ${\dot{\bm{\mathrm{P}}}}^{(2)}$ into $P$ overlapping patches respectively;
10Divide each quaternion representation of ${\{\dot{\bm{\mathrm{I}}}^{(i)}\}}_{i=1}^{n}$ and ${\{\dot{\bm{\mathrm{D}}}^{(i)}\}}_{i=1}^{n}$ into $P$ overlapping patches respectively;
11Compute $QWSSIM$ values at the patch level using Eq. (17);
12Obtain the refined focus map and fused patches at the patch level using Eqs. (19) and (20);
Aggregate the patches of ${\{\dot{\bm{\mathrm{F}}}_{j}\}}_{j=1}^{P}$ to obtain the final result $\bm{\mathrm{F}}$;
Algorithm 2 Quaternion Multi-focus Color Image Fusion

### v2

@@ III-B1 Quaternion focal element decomposition
1 Initialize ${\{\dot{\bm{\mathrm{J}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$${\{\dot{\bm{\mathrm{Z}}}{{}_{k}^{0}}\}}{{}_{k=1}^{K}}$, ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}^{0}}\}}{{}_{k=1}^{K}}$${\{\dot{\bm{\mathrm{Y}}}{{}_{2,k}^{0}}\}}{{}_{k=1}^{K}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}^{0}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}^{0}}$, $\dot{\bm{\mathrm{B}}}{{}^{0}}$, $\dot{\bm{\mathrm{D}}}{{}^{0}}$, $\dot{\bm{\mathrm{E}}}{{}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}^{0}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}^{0}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{5}^{0}}$;
@@ 9  Fix other variables and solve $\dot{\bm{\mathrm{D}}}{{}^{t}}$ using Lemma 2;
[…]
@@
[…]
@@ 20 $t$ $\longleftarrow$ $t+1$;
Algorithm 1 Quaternion focal element decomposition (QFED)
III-B2 Patch-wise dual-scale focus map generation
As illustrated in Fig. 2, we obtain the corresponding detail layers $\dot{\bm{\mathrm{D}}}{{}_{1}}$, $\dot{\bm{\mathrm{D}}}{{}_{2}}$ and coefficient matrices $\dot{\bm{\mathrm{Z}}}{{}_{1}}$, $\dot{\bm{\mathrm{Z}}}{{}_{2}}$ from the QFED process for given input quaternion representations $\dot{\bm{\mathrm{I}}}{{}_{1}}$, $\dot{\bm{\mathrm{I}}}{{}_{2}}$. To fully leverage the complementary information in $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$, we propose a patch-wise dual-scale focus map generation strategy to generate the base-scale and detail-scale focus maps at the patch level.
Detail Amplification. The detail-scale layer for each pixel location is amplified within a sliding window to strengthen the fine-grained details:
$\dot{\bm{\mathrm{D}}}{{}_{s}}(x,y)=\sum_{s=-r}^{r}\sum_{t=-r}^{r}\dot{\bm{\mathrm{D}}}(x+s,y+t),$ (13)
[…]
Dual-scale focus measure. The dual-scale focus measures used to compute patch-wise focus levels are defined as:
$l_{B}={\|\nabla_{1}\dot{\bm{\mathrm{d}}}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{d}}}\|}_{1}+\theta{\|\dot{\bm{\mathrm{z}}}\|}_{2},$ (14)
$l_{D}=\phi({\|\nabla_{1}\dot{\bm{\mathrm{d}}}_{s}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{d}}}_{s}\|}_{1}),$ (15)
where $l_{B}$ and $l_{D}$ represent patch-wise base-scale and detail-scale focus levels of input quaternion representation $\dot{\bm{\mathrm{I}}}$; $\dot{\bm{\mathrm{d}}}$ and $\dot{\bm{\mathrm{z}}}$ denote the local patch and corresponding column vector of $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$ respectively; $\dot{\bm{\mathrm{d}}}_{s}$ represents the local patch of $\dot{\bm{\mathrm{D}}}_{s}$ computed by Eq. (13). Parameter $\theta$ is set to 1 by default. $\phi(\cdot)$ is a Laplacian-based enhancement function defined as $\phi(x)=1-e^{-x/\gamma}$ where $\gamma=0.2$ by default. The detail-scale focus measure captures the magnitude of two-directional gradients that effectively reflect focus levels in color images. The base-scale focus measure jointly considers gradient magnitudes and coefficient matrix energy, thereby improving spatial coherence and mitigating local artifacts.
To compare base-scale and detail-scale focus levels of the local patches in $\dot{\bm{\mathrm{I}}}{{}_{1}}$ and $\dot{\bm{\mathrm{I}}}{{}_{2}}$, the patch-wise dual-scale focus maps $\bm{\mathrm{m}}{{}_{B}}$ and $\bm{\mathrm{m}}{{}_{D}}$ are computed as:
$\bm{\mathrm{m}}{{}_{B}}=\begin{cases}0,&{\rm if}\ l_{B,1}>l_{B,2}\\ 1,&{\rm if}\ l_{B,1}\leq l_{B,2}\end{cases}$ (16)
$\bm{\mathrm{m}}{{}_{D}}=\begin{cases}0,&{\rm if}\ l_{D,1}>l_{D,2}\\ 1,&{\rm if}\ l_{D,1}\leq l_{D,2}\end{cases}$ (17)
where $l_{B,1}$ and $l_{B,2}$ are computed using Eq. (14); $l_{D,1}$ and $l_{D,2}$ are computed using Eq. (15); $\bm{\mathrm{m}}{{}_{B}}$ and $\bm{\mathrm{m}}{{}_{D}}$ denote patch-wise base-scale and detail-scale focus maps respectively. Aggregating these decisions across all patches yields the global base-scale and detail-scale focus maps $\bm{\mathrm{M}}{{}_{B}}$ and $\bm{\mathrm{M}}{{}_{D}}$.
Figure 3: Our quaternion base-detail fusion strategy.
@@ III-C Quaternion base-detail fusion strategy
Our quaternion base-detail fusion (QBDF) strategy aims to achieve high-quality multi-focus color image fusion that operates at base and detail scales. The overall fusion process is illustrated in Fig. 3. Given the input quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$ $\in\mathbbm{H}^{M\times N}$ and $\dot{\bm{\mathrm{I}}}_{2}$ $\in\mathbbm{H}^{M\times N}$ and corresponding dual-scale focus maps $\bm{\mathrm{M}}_{B}$ and $\bm{\mathrm{M}}_{D}$, the QBDF strategy performs dual-scale image fusion.
The patch-wise selection results that are derived from the focus maps are used to construct the fused images at each scale. Specifically, the base-scale fused result ${\dot{\bm{\mathrm{F}}}}_{1}$ and detail-scale fused result ${\dot{\bm{\mathrm{F}}}}_{2}$ are computed using the following rules:
${\dot{\bm{\mathrm{F}}}}_{1}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{I}}}}_{1},&&{\rm if}\ \bm{\mathrm{M}}_{B}=0\\ &{\dot{\bm{\mathrm{I}}}}_{2},&&{\rm if}\ \bm{\mathrm{M}}_{B}=1\end{aligned}\right$. (18)
${\dot{\bm{\mathrm{F}}}}_{2}=\left\{\begin{aligned} &{\dot{\bm{\mathrm{I}}}}_{1},&&{\rm if}\ \bm{\mathrm{M}}_{D}=0\\ &{\dot{\bm{\mathrm{I}}}}_{2},&&{\rm if}\ \bm{\mathrm{M}}_{D}=1\end{aligned}\right$. (19)
This fusion rule ensures that the source image with the most focused content is selected and integrated at detail or base scales for each patch location.
@@ III-D Quaternion structural similarity refinement strategy
To further enhance the quality of the final fused image, we propose a quaternion structural similarity refinement (QSSR) strategy. It refines the local patches of base-scale and detail-scale fusion results $\dot{\bm{\mathrm{F}}}_{1}$ and $\dot{\bm{\mathrm{F}}}_{2}$ using a patch-wise selection rule. The assumption is that the optimal fused patch should exhibit the highest similarity with the corresponding region in the source image that contains the most focused content. We first propose a weighted quaternion structure similarity ($WQ_{SSIM}$) measure.
Weighted quaternion structural similarity measure. For each spatial location, let $\dot{\bm{\mathrm{f}}}$ denote a local patch of a base-scale or detail-scale fusion result $\dot{\bm{\mathrm{F}}}_{2}$ or $\dot{\bm{\mathrm{F}}}_{2}$. $\dot{\bm{\mathrm{p}}}_{1}$ and $\dot{\bm{\mathrm{p}}}_{2}$ represent the corresponding local patches extracted from $\dot{\bm{\mathrm{I}}}_{1}$ and $\dot{\bm{\mathrm{I}}}_{2}$ at the same spatial location. The $WQ_{SSIM}$ value is computed as follows:
$WQ_{SSIM}({\dot{\bm{\mathrm{f}}}})={\tau}_{1}Q_{SSIM}(\dot{\bm{\mathrm{f}}},\dot{\bm{\mathrm{p}}}_{1})+{\tau}_{2}Q_{SSIM}(\dot{\bm{\mathrm{f}}},\dot{\bm{\mathrm{p}}}_{2})$ (20)
where quaternion structural similarity ($Q_{SSIM}$) is defined in Definition 5 in Section II; ${\tau}_{1}$ and ${\tau}_{2}$ are the adaptive weights assigned to $\dot{\bm{\mathrm{p}}}_{1}$ and $\dot{\bm{\mathrm{p}}}_{2}$ respectively based on its detail-scale focus level. They are defined below:
${\tau}_{1}=\frac{l_{D,1}}{l_{D,1}+l_{D,2}+\epsilon},\quad{\tau}_{2}=1-{\tau}_{1}$ (21)
where $\epsilon$ is a small positive constant.
$WQ_{SSIM}$ guides the selection of the optimal patch at each spatial location. The final fused patch $\dot{\bm{\mathrm{p}}}$ at each location is selected according to:
$\dot{\bm{\mathrm{p}}}=\left\{\begin{aligned} &\dot{\bm{\mathrm{f}}}_{1},&{\rm if}\ WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{1})>WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{2})\\ &\dot{\bm{\mathrm{f}}}_{2},&{\rm if}\ WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{1})\leq WQ_{SSIM}({\dot{\bm{\mathrm{f}}}}_{2})\end{aligned}\right$. (22)
This strategy aims to effectively suppress artifacts and corrects misclassified regions in the fused outputs. Aggregating all selected patches according to Eq. (22) yields the quaternion fused result $\dot{\bm{\mathrm{F}}}$. Convert $\dot{\bm{\mathrm{F}}}$ back to the real domain to obtain the fused color image $\bm{\mathrm{F}}$.